OpenStudy (anonymous):

Find the line that passes through the point (-2,3) and is parallel to the line y=2x-3. What information are you given? What equation can you see?

OpenStudy (radar):

You are given the slope of the line which you are to find. Where you say? It is given that it is parallel to the line y=2x-3, parallel lines have equal slope so the slope of the line you are to find is 2. (the coefficient of x).

OpenStudy (radar):

Secondly you are given a point (-2,3) which the line you are to find passes through. What else could you want?

OpenStudy (radar):

You can now use the equation: m=(y-3)/(x-(-2)) and obtain the equation of the desired line you are to find.

OpenStudy (radar):

Plug in 2 for m and solve.

OpenStudy (radar):

I get y=2x+7 What did you get?

OpenStudy (anonymous):

y=2x-5 i guess thats wrong

OpenStudy (radar):

Let me double check my work. 2 (y -3) -- = ----- cross multiplying 1 (x +2) 2x+4 =y - 3 y=2x + 7 appears to be correct

OpenStudy (radar):

How did you come up with the -5

OpenStudy (radar):

*I think I see what you did, you didn't treat the double negative correctly.

OpenStudy (radar):

x - (-2) becomes x + 2

OpenStudy (anonymous):

y-1=m(x-x1) y- -1=2(x-2) y+1=2(x-2) y+1=2x-4 -1 -1 y=2x-5

OpenStudy (radar):

it is x - (-2) That becomes x+2

OpenStudy (radar):

Do you see that? The P is (-2,3) x=-2, y=3

OpenStudy (radar):

All other operations were correct, you're almost there.

OpenStudy (anonymous):

i see, thanks i got it now

OpenStudy (radar):

Great, good luck with your studies.

OpenStudy (anonymous):

thank you